# On the isotopies of tangles in periodic 3-manifolds using finite covers

### v1

@@ 3.2. Links in the thickened torus
For $i=0,1$, let $d_{i}=\gcd(p_{i},q_{i})$, $p_{i}^{\prime}=p_{i}/d_{i}$, and $q_{i}^{\prime}=q_{i}/d_{i}$. Since the number of component of the link $T_{2}(p_{i},q_{i})$ is $d_{i}$, we have $d_{0}=d_{1}$. Since any component $T_{2}(p_{i},q_{i})$ represents ±$\pm(p_{i}^{\prime},q_{i}^{\prime})\in\mathbb{Z}^{2}=\pi_{1}(T^{2}\times I)$, we have $p_{0}^{\prime}=p_{1}^{\prime}$ and $q_{0}^{\prime}=q_{1}^{\prime}$. Hence $p_{0}=p_{1}$ and $q_{0}=q_{1}$. ∎
@@
In the first case, we say that $T$ is a layering torus. In the last case, the other region bounded by $T$ is called a knotted hole ball.
@@
Fix a Seifert fibration of $T^{2}\times I$, which has no singular fibers. If $T$ is incompressible, then $T$ is isotopic to be vertical (a union of regular fibers) or horizontal (transverse to all fibers) in the Seifert fibration by [25, Proposition 1.11]. Since $T$ is closed, $T$ is vertical. Hence $T$ is parallel to the boundary of $T^{2}\times I$. If $T$ is compressible, then $T$ bounds a solid torus or a knotted hole ball in the same manner as Lemma 3.3. ∎
@@
The classification in [6, Proposition 3.2] implies the following classification.
@@ 3.3. Links in the 3-torus
For $i=0,1$, let $d_{i}=\gcd(p_{i},q_{i},r_{i})$, $p_{i}^{\prime}=p_{i}/d_{i}$, $q_{i}^{\prime}=q_{i}/d_{i}$, and $r_{i}^{\prime}=r_{i}/d_{i}$. Since the number of component of the link $T_{3}(p_{i},q_{i},r_{i})$ is $d_{i}$, we have $d_{0}=d_{1}$. Since any component $T_{3}(p_{i},q_{i},r_{i})$ represents ±$\pm(p_{i}^{\prime},q_{i}^{\prime},r_{i}^{\prime})\in\mathbb{Z}^{3}=\pi_{1}(X)$, we have $p_{0}^{\prime}=p_{1}^{\prime}$, $q_{0}^{\prime}=q_{1}^{\prime}$, and $r_{0}^{\prime}=r_{1}^{\prime}$. Hence $p_{0}=p_{1}$, $q_{0}=q_{1}$, and $r_{0}=r_{1}$. ∎
@@
Hence the meridian of any component of $L$ is not a fiber in the Seifert fibration on $T^{3}\setminus\!\!\setminus L$. In this case, the Seifert fibration on $T^{3}\setminus\!\!\setminus L$ extends to $T^{3}$. Any Seifert fibration of the 3-torus is isomorphic to the product fibering of $M(1,0;)$ by [25, Theorem 2.3]. Moreover, $L$ consists of regular fibers. Therefore $L$ is isotopic to $T_{3}(p,q,r)$. ∎
@@
In the first case, we say that $T$ is a layering torus. In the last case, the other region bounded by $T$ is called a knotted hole ball.
@@
Fix a Seifert fibration of $T^{3}$, which has no singular fibers. If $T$ is incompressible, then $T$ is isotopic to be vertical or horizontal in the Seifert fibration by [25, Proposition 1.11]. In both cases, $T^{3}\setminus\!\!\setminus T$ is a thickened torus. If $T$ is compressible, then $T$ bounds a solid torus or a knotted hole ball in the same manner as Lemma 3.3. ∎
@@ 4. Cover-isotopy of links
We first show that cover-isotopy implies isotopy for hyperbolic links in the solid torus, the thickened torus, or the 3-torus. We need to compare group actions on a hyperbolic 3-manifold. For a group $G$ and a topological space $M$, let $\alpha\colon G\to\operatorname{Aut}(M)$ be a homomorphism, called a (left) action of $G$ on $M$, where $\operatorname{Aut}(M)$ is the group consisting of the self-homeomorphisms of $M$. An action is free if $\alpha(g)$ has no fixed points for each non-trivial element $g\in G$. A homeomorphism $f\colon M\to M^{\prime}$ induces the action $f_{*}(\alpha)$ on $M^{\prime}$ defined by $f_{*}(\alpha)(g)=f\circ\alpha(g)\circ f^{-1}\in\operatorname{Aut}(M^{\prime})$ for $g\in G$. An isometric action of $G$ on a hyperbolic 3-manifold $M$ is a homomorphism $\alpha\colon G\to\operatorname{Isom}(M)$, where $\operatorname{Isom}(M)$ is the group consisting of the self-isometries of $M$. An isometric action on a hyperbolic 3-manifold induces an action on the cusps which is isometric with respect to the Euclidean structure. If this action is free, the group acts to the tori by translations and interchanging components.
@@
which is an analogue of the rotational number of a self-homeomorphism of $S^{1}$. If $\tilde{f}\colon\mathbb{R}^{m}\to\mathbb{R}^{m}$ is a lift of $f\in\operatorname{Aut}_{0}(T^{m})$, then $\rho(\tilde{f}\circ\tilde{h}\circ\tilde{f}^{-1})=\rho(\tilde{h})$.
@@
If $h\colon T^{m}\to T^{m}$ is an isometry isotopic to the identity, then $h\in\operatorname{Aut}_{0}(T^{m})$ is a translation by an element $v\in\mathbb{R}^{m}/\mathbb{Z}^{m}$. The rotation set $\rho(\tilde{h})$ is determined as a single point $\tilde{v}\in\mathbb{R}^{m}$ modulo $\mathbb{Z}^{m}$ that is a lift of $v$. Since there are lifts $\tilde{h}_{0},\tilde{h}_{1},\tilde{f}\colon\mathbb{R}^{m}\to\mathbb{R}^{m}$ of $h_{0},h_{1},f\colon T^{m}\to T^{m}$ such that $\tilde{f}\circ\tilde{h}_{0}\circ\tilde{f}^{-1}=\tilde{h}_{1}$, we have $\rho(\tilde{h}_{0})\equiv\rho(\tilde{h}_{1})$ modulo $\mathbb{Z}^{m}$. Hence the translations $h_{0}$ and $h_{1}$ coincide. ∎
We state a sufficient condition for cover-isotopy of hyperbolic links to imply isotopy in a general setting. Suppose that $X$ is a 3-manifold with (possibly empty) boundary consisting of tori. Let $L_{0}$ and $L_{1}$ be hyperbolic links in $X$. Let $P\colon\widetilde{X}\to X$ be a finite regular covering map. Suppose that the lifts $\widetilde{L_{0}}=P^{-1}(L_{0})$ and $\widetilde{L_{1}}=P^{-1}(L_{1})$ are isotopic in $\widetilde{X}$. Then the interior of the complement $M_{i}=X\setminus\!\!\setminus L_{i}$ admits a hyperbolic metric for $i=0,1$. Its cover $\widetilde{M_{i}}=\widetilde{X}\setminus\!\!\setminus\widetilde{L_{i}}$ is endowed with the lifted hyperbolic metric. The isotopy between $\widetilde{L_{0}}$ and $\widetilde{L_{1}}$ induces a homeomorphism $f\colon\widetilde{M_{0}}\to\widetilde{M_{1}}$. The Mostow rigidity implies that $f$ is isotopic to an isometry $\tilde{\iota}\colon\widetilde{M_{0}}\to\widetilde{M_{1}}$. The finite group $G=\pi_{1}(X)/\pi_{1}(\widetilde{X})$ acts on $\widetilde{X}$ by the deck transformations. This action induces free isometric actions of $G$ on the hyperbolic 3-manifolds $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$, denoted by $\alpha_{0}$ and $\alpha_{1}$.
@@
The homeomorphism $\iota\colon(X,L_{0})\to(X,L_{1})$ lifts to a homeomorphism $\tilde{\iota}\colon(\widetilde{X},\widetilde{L_{0}})\to(\widetilde{X},\widetilde{L_{1}})$ isotopic to the identity of $\widetilde{X}$. If $\iota$ is isotopic to the identity of $X$, then $L_{0}$ and $L_{1}$ are isotopic in $X$. ∎
@@
Recall that $\operatorname{Aut}(X)$ is the group consisting of self-homeomorphisms of $X$, and $\operatorname{Aut}_{0}(X)$ is the subgroup of $\operatorname{Aut}(X)$ consisting of self-homeomorphisms of $X$ isotopic to the identity. The mapping class group of $X$ is defined as the quotient group $\pi_{0}(\operatorname{Aut}(X))=\operatorname{Aut}(X)/\operatorname{Aut}_{0}(X)$. The action on the fundamental group induces the natural homomorphism from $\pi_{0}(\operatorname{Aut}(X))$ to the outer automorphism group $\operatorname{Out}(\pi_{1}(X))$. Moreover, $\operatorname{Out}(\mathbb{Z}^{n})\cong\mathrm{GL}(n,\mathbb{Z})$. The following isomorphisms are known to hold:
@@
Since $\pi_{1}(X)$ is abelian, the isomorphism $f_{*}\colon\pi_{1}(X)\to\pi_{1}(X)$ does not depend on the choice of a base point. A self-homeomorphism $f\in\operatorname{Aut}(X)$ lifts to $\tilde{f}\in\operatorname{Aut}(\widetilde{X})$ if and only if the two subgroups $f_{*}\circ P_{*}(\pi_{1}(\widetilde{X}))$ and $P_{*}(\pi_{1}(\widetilde{X}))$ of $\pi_{1}(X)$ coincide. In other words, the subgroup $P_{*}(\pi_{1}(\widetilde{X}))$ of $\pi_{1}(X)$ is invariant under $f_{*}$. Then $\tilde{f}_{*}\colon\pi_{1}(\widetilde{X})\to\pi_{1}(\widetilde{X})$ is the restriction of $f_{*}$.
Suppose that $\tilde{f}\in\operatorname{Aut}_{0}(\widetilde{X})$. Then $\tilde{f}_{*}\colon\pi_{1}(\widetilde{X})\to\pi_{1}(\widetilde{X})$ is the identity. Hence $f_{*}\colon\pi_{1}(X)\to\pi_{1}(X)$ is also the identity. In other words, $[f]\in\pi_{0}(\operatorname{Aut}(X))$ is mapped to the identity in $\operatorname{Out}(\pi_{1}(X))$. Since $\tilde{f}$ is orientation-preserving, so is $f$. Hence $f\in\operatorname{Aut}_{0}(X)$ in the cases that $X$ is $T^{2}\times I$ or $T^{3}$. In the case that $X$ is $S^{1}\times D^{2}$, we need to pay attention to the $\mathbb{Z}$-subgroup of $\pi_{0}(\operatorname{Aut}(S^{1}\times D^{2}))$. If its element for $[f]\in\pi_{0}(\operatorname{Aut}(X))$ is $a$, then that for $[\tilde{f}]\in\pi_{0}(\operatorname{Aut}(\widetilde{X}))$ is the product of $a$ and the degree of the covering map $P\colon\widetilde{X}\to X$, which is equal to zero. Hence $a=0$. Therefore $f\in\operatorname{Aut}_{0}(X)$. ∎
Note that if $X$ is $S^{1}\times S^{2}$, the conclusion in Lemma 4.3 does not hold. Indeed, $\pi_{0}(\operatorname{Aut}(S^{1}\times S^{2}))\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Its $\mathbb{Z}/2\mathbb{Z}$-subgroup instead of the $\mathbb{Z}$-subgroup of $\pi_{0}(\operatorname{Aut}(S^{1}\times D^{2}))$ is due to $\pi_{1}(\mathrm{SO}(3))\cong\mathbb{Z}/2\mathbb{Z}$ (cf. the belt trick). Hence a self-homeomorphism of $S^{1}\times S^{2}$ corresponding to $(0,1,0)\in\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ lifts to a self homeomorphism of the double cover isotopic to the identity.
@@
Since the fundamental group $\pi_{1}(X)$ is abelian, the covering map $P$ is regular. We already have an isometry $\tilde{\iota}\colon\widetilde{M_{0}}\to\widetilde{M_{1}}$ and a free isometric action $\alpha_{i}$ of the finite group $G=\pi_{1}(X)/\pi_{1}(\widetilde{X})$ on the hyperbolic 3-manifold $\widetilde{M_{i}}$ for each $i=0,1$. By Lemmas 4.2 and 4.3, it is sufficient to show that $\tilde{\iota}_{*}(\alpha_{0})=\alpha_{1}$.
@@
We next suppose that $X$ is $T^{3}$. We consider the two subgroups $G_{0}=\tilde{\iota}_{*}(\alpha_{0})(G)$ and $G_{1}=\alpha_{1}(G)$ of the finite group $\operatorname{Isom}(\widetilde{M_{1}})$. Let $\overline{G}$ denote the subgroup of $\operatorname{Isom}(\widetilde{M_{1}})$ generated by $G_{0}$ and $G_{1}$. Since the elements of $G_{0}$ and $G_{1}$ preserves the meridians, so do the elements of $\overline{G}$. Hence the action of the finite group $\overline{G}$ on $\widetilde{M_{1}}$ can be extended to a smooth action on $\widetilde{X}$. Note that this action might not be free. The 3-manifold or 3-orbifold $\widetilde{X}/\overline{G}$ is a quotient of a 3-torus. The geometrization of 3-manifolds and 3-orbifolds [4, 11, 61] implies that $\widetilde{X}/\overline{G}$ admits a Euclidean structure. Hence $\widetilde{X}$ admits a Euclidean structure such that the action of $\overline{G}$ on $\widetilde{X}$ is isometric. The actions $\tilde{\iota}_{*}(\alpha_{0})$ and $\alpha_{1}$ of $G$ on the Euclidean torus $\widetilde{X}$ are isometric. Since they are conjugate by a self-homeomorphism of $\widetilde{X}$ isotopic to the identity, we have $\tilde{\iota}_{*}(\alpha_{0})=\alpha_{1}$ by Lemma 4.1. ∎
@@
We give an example to note. Let $L_{0}=T_{0}(2,0)$ and $L_{1}=T_{0}(2,1)$. Since the numbers of components are different, there is no self-homeomorphism of $X$ that maps $L_{0}$ to $L_{1}$. Suppose that $\widetilde{X}$ is the double cover of $X$. Then there is a self-homeomorphism of $\widetilde{X}$ that maps $\widetilde{L_{0}}=T_{0}(2,0)$ to $\widetilde{L_{1}}=T_{0}(2,2)$ as shown in Figure 5. However, they are not isotopic in fixed $\widetilde{X}$. Hence this is not a counterexample of Theorem 4.5. There are such examples of hyperbolic links as shown in Figure 6.
@@
Proposition 3.2 implies that $L_{0}$ is isotopic to $T_{l_{0}}(p_{0},q_{0})$, and $L_{1}$ is isotopic to $T_{l_{1}}(p_{1},q_{1})$. We may assume that $p_{0},p_{1}\geq 0$, and $p_{i}$ does not divide $q_{i}$ if $l_{i}=0$. Then $\widetilde{L_{0}}$ is isotopic to $T_{l_{0}}(p_{0},nq_{0})$, and $\widetilde{L_{1}}$ is isotopic to $T_{l_{1}}(p_{1},nq_{1})$. Since $\widetilde{L_{0}}$ and $\widetilde{L_{1}}$ are isotopic, we have $l_{0}=l_{1}$, $p_{0}=p_{1}$, and $q_{0}=q_{1}$ by Lemma 3.1. Hence $L_{0}$ and $L_{1}$ are isotopic. ∎
@@
We consider the JSJ decomposition of the irreducible 3-manifolds $X\setminus\!\!\setminus L_{0}$ and $X\setminus\!\!\setminus L_{1}$. The JSJ decomposition of $\widetilde{X}\setminus\!\!\setminus\widetilde{L_{0}}\simeq\widetilde{X}\setminus\!\!\setminus\widetilde{L_{1}}$ is their lifts by Lemma 2.11. We show the assertion by induction of the number $N$ of JSJ pieces of $\widetilde{X}\setminus\!\!\setminus\widetilde{L_{0}}$. We first suppose that $\widetilde{X}\setminus\!\!\setminus\widetilde{L_{0}}(\simeq\widetilde{X}\setminus\!\!\setminus\widetilde{L_{1}})$ has no essential tori. Then the link $\widetilde{L_{0}}$ is either Seifert fibered or hyperbolic, and the links $L_{0}$ and $L_{1}$ are in the same class. In this case, the assertion holds by Theorem 4.4 and Lemma 4.6.
Suppose that $N>1$ and the assertion holds for any degree $n$ of the covering map if the number of JSJ pieces is less than $N$. Let $M_{0}$ and $M_{1}$ respectively denote the outermost JSJ pieces of $X\setminus\!\!\setminus L_{0}$ and $X\setminus\!\!\setminus L_{1}$. Their preimages $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ in $\widetilde{X}$ are isotopic in $\widetilde{X}$. The 3-manifolds $M_{0}$ and $M_{1}$ can be reembedded to the complements of links in $X$ by Lemma 3.6. Here a knotted hole ball for $M_{i}$ is lifted to its disjoint $n$ copies in $\widetilde{X}$. The 3-manifolds $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ are equivariantly reembedded. Since $M_{0}$ and $M_{1}$ are either Seifert fibered or hyperbolic, $M_{0}$ and $M_{1}$ are isotopic in $X$ by Theorem 4.4 and Lemma 4.6. Note that if $M_{i}$ is Seifert fibered and has a boundary component bounding a knotted hole ball, then $M_{i}$ is reembedded to $X\setminus\!\!\setminus T_{1}(0,q)$.
@@
Proposition 3.9 implies that $L_{0}$ is isotopic to $T_{2}(p_{0},q_{0})$, and $L_{1}$ is isotopic to $T_{2}(p_{1},q_{1})$. We may assume that $p_{0},p_{1}\geq 0$. We take $(\bm{a},\bm{b})$ as a basis for $\pi_{1}(\widetilde{X})$. Then $\widetilde{L_{i}}$ for $i=0,1$ is isotopic to $T_{2}(\tilde{p}_{i},\tilde{q}_{i})$, where
@@
Let $M_{0}$ and $M_{1}$ respectively denote the outermost JSJ pieces of $X\setminus\!\!\setminus L_{0}$ and $X\setminus\!\!\setminus L_{1}$ as in Lemma 3.12. Their preimages $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ in $\widetilde{X}$ are isotopic in $\widetilde{X}$. In the same manner as Theorem 4.5, $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ are equivariantly reembedded in the thickened torus. Since $M_{0}$ and $M_{1}$ are either Seifert fibered or hyperbolic, $M_{0}$ and $M_{1}$ are isotopic in $X$ by Theorem 4.4 and Lemma 4.8.
@@
Proposition 3.15 implies that $L_{0}$ is isotopic to $T_{3}(p_{0},q_{0},r_{0})$, and $L_{1}$ is isotopic to $T_{3}(p_{1},q_{1},r_{1})$. We may assume that $p_{0},p_{1}\geq 0$. We take $(\bm{a},\bm{b},\bm{c})$ as a basis for $\pi_{1}(\widetilde{X})$. Then $\widetilde{L_{i}}$ for $i=0,1$ is isotopic to $T_{3}(\tilde{p}_{i},\tilde{q}_{i},\tilde{r}_{i})$, where

### v2

@@ 3.2. Links in the thickened torus
For $i=0,1$, let $d_{i}=\gcd(p_{i},q_{i})$, $p_{i}^{\prime}=p_{i}/d_{i}$, and $q_{i}^{\prime}=q_{i}/d_{i}$. Since the number of component of the link $T_{2}(p_{i},q_{i})$ is $d_{i}$, we have $d_{0}=d_{1}$. Since any component $T_{2}(p_{i},q_{i})$ represents $\pm(p_{i}^{\prime},q_{i}^{\prime})\in\mathbb{Z}^{2}=\pi_{1}(T^{2}\times I)$, we have $p_{0}^{\prime}=p_{1}^{\prime}$ and $q_{0}^{\prime}=q_{1}^{\prime}$. Hence $p_{0}=p_{1}$ and $q_{0}=q_{1}$. ∎
@@
In the first case, we say that $T$ is a layering torus. In the last case, $T$ bounds a solid torus or a knotted hole ball.
@@
Fix a Seifert fibration of $T^{2}\times I$, which has no singular fibers. If $T$ is incompressible, then $T$ is isotopic to be vertical (a union of regular fibers) or horizontal (transverse to all fibers) in the Seifert fibration by [29, Proposition 1.11]. Since $T$ is closed, $T$ is vertical. Hence $T$ is parallel to the boundary of $T^{2}\times I$. If $T$ is compressible, then $T$ bounds a solid torus or a knotted hole ball in the same manner as Lemma 3.3. ∎
@@
The classification in [7, Proposition 3.2] implies the following classification.
@@ 3.3. Links in the 3-torus
For $i=0,1$, let $d_{i}=\gcd(p_{i},q_{i},r_{i})$, $p_{i}^{\prime}=p_{i}/d_{i}$, $q_{i}^{\prime}=q_{i}/d_{i}$, and $r_{i}^{\prime}=r_{i}/d_{i}$. Since the number of component of the link $T_{3}(p_{i},q_{i},r_{i})$ is $d_{i}$, we have $d_{0}=d_{1}$. Since any component $T_{3}(p_{i},q_{i},r_{i})$ represents $\pm(p_{i}^{\prime},q_{i}^{\prime},r_{i}^{\prime})\in\mathbb{Z}^{3}=\pi_{1}(X)$, we have $p_{0}^{\prime}=p_{1}^{\prime}$, $q_{0}^{\prime}=q_{1}^{\prime}$, and $r_{0}^{\prime}=r_{1}^{\prime}$. Hence $p_{0}=p_{1}$, $q_{0}=q_{1}$, and $r_{0}=r_{1}$. ∎
@@
Hence the meridian of any component of $L$ is not a fiber in the Seifert fibration on $T^{3}\setminus\!\!\setminus L$. In this case, the Seifert fibration on $T^{3}\setminus\!\!\setminus L$ extends to $T^{3}$. Any Seifert fibration of the 3-torus is isomorphic to the product fibering of $M(1,0;)$ by [29, Theorem 2.3]. Moreover, $L$ consists of regular fibers. Therefore $L$ is isotopic to $T_{3}(p,q,r)$. ∎
@@
In the first case, we say that $T$ is a layering torus. In the last case, $T$ bounds a solid torus or a knotted hole ball.
@@
Fix a Seifert fibration of $T^{3}$, which has no singular fibers. If $T$ is incompressible, then $T$ is isotopic to be vertical or horizontal in the Seifert fibration by [29, Proposition 1.11]. In both cases, $T^{3}\setminus\!\!\setminus T$ is a thickened torus. If $T$ is compressible, then $T$ bounds a solid torus or a knotted hole ball in the same manner as Lemma 3.3. ∎
@@ 4. Cover-isotopy of links
We first show that cover-isotopy implies isotopy for hyperbolic links in the solid torus, the thickened torus, or the 3-torus. We need to compare group actions on a hyperbolic 3-manifold. For a group $G$ and a topological space $M$, let $\alpha\colon G\to\operatorname{Aut}(M)$ be a homomorphism, called a (left) action of $G$ on $M$, where $\operatorname{Aut}(M)$ is the group consisting of the self-homeomorphisms of $M$. An action is free if $\alpha(g)$ has no fixed points for each non-trivial element $g\in G$. A homeomorphism $f\colon M\to M^{\prime}$ induces the action $f_{*}(\alpha)$ on $M^{\prime}$ defined by $f_{*}(\alpha)(g)=f\circ\alpha(g)\circ f^{-1}\in\operatorname{Aut}(M^{\prime})$ for $g\in G$. An isometric action of $G$ on a hyperbolic 3-manifold $M$ is a homomorphism $\alpha\colon G\to\operatorname{Isom}(M)$, where $\operatorname{Isom}(M)$ is the group consisting of the self-isometries of $M$. An isometric action on a hyperbolic 3-manifold induces an action on the cusps which is isometric with respect to the Euclidean structure. If this action is free, the group acts on the tori by translations and interchanging components.
@@
[…]
@@
[…]
@@
[…]
@@
Recall that $\operatorname{Aut}(X)$ is the group consisting of self-homeomorphisms of $X$, and $\operatorname{Aut}_{0}(X)$ is the subgroup of $\operatorname{Aut}(X)$ consisting of self-homeomorphisms of $X$ isotopic to the identity. The mapping class group of $X$ is defined as the quotient group $\pi_{0}(\operatorname{Aut}(X))=\operatorname{Aut}(X)/\operatorname{Aut}_{0}(X)$. The action on the fundamental group induces the natural homomorphism from $\pi_{0}(\operatorname{Aut}(X))$ to the outer automorphism group $\operatorname{Out}(\pi_{1}(X))$. Moreover, $\operatorname{Out}(\mathbb{Z}^{n})\cong\mathrm{GL}(n,\mathbb{Z})$. Due to Hatcher [28], the following isomorphisms are known to hold:
@@
[…]
Note that if $X$ is $S^{1}\times S^{2}$, the conclusion in Lemma 4.3 does not hold. Indeed, $\pi_{0}(\operatorname{Aut}(S^{1}\times S^{2}))\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. Its $\mathbb{Z}/2\mathbb{Z}$-subgroup instead of the $\mathbb{Z}$-subgroup of $\pi_{0}(\operatorname{Aut}(S^{1}\times D^{2}))$ is due to $\pi_{1}(\mathrm{SO}(3))\cong\mathbb{Z}/2\mathbb{Z}$ (cf. the belt trick). Hence a self-homeomorphism of $S^{1}\times S^{2}$ corresponding to $(0,1,0)\in\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ lifts to a self-homeomorphism of the double cover isotopic to the identity.
@@
[…]
@@
[…]
@@
We give an example to note. Let $L_{0}=T_{0}(2,0)$ and $L_{1}=T_{0}(2,1)$. Since the numbers of their components are different, there is no self-homeomorphism of $X$ that maps $L_{0}$ to $L_{1}$. Suppose that $\widetilde{X}$ is the double cover of $X$. Then there is a self-homeomorphism of $\widetilde{X}$ that maps $\widetilde{L_{0}}=T_{0}(2,0)$ to $\widetilde{L_{1}}=T_{0}(2,2)$ as shown in Figure 5. However, they are not isotopic in fixed $\widetilde{X}$. Hence this is not a counterexample of Theorem 4.5. There are such examples of hyperbolic links as shown in Figure 6.
@@
Proposition 3.2 implies that $L_{0}$ is isotopic to $T_{l_{0}}(p_{0},q_{0})$, and $L_{1}$ is isotopic to $T_{l_{1}}(p_{1},q_{1})$. We may assume that $p_{0},p_{1}\geq 0$, and $p_{i}$ does not divide $q_{i}$ if $l_{i}=0$. Then $\widetilde{L_{0}}$ is isotopic to $T_{l_{0}}(p_{0},nq_{0})$, and $\widetilde{L_{1}}$ is isotopic to $T_{l_{1}}(p_{1},nq_{1})$. Hence $T_{l_{0}}(p_{0},nq_{0})$ and $T_{l_{1}}(p_{1},nq_{1})$ are isotopic. If $l_{0}=0$ and $p_{0}$ divides $nq_{0}$, then $T_{0}(p_{0},nq_{0})$ is isotopic to $T_{1}(p_{0}-1,n(p_{0}-1)q_{0}/p_{0})$. In this case, since $p_{0}$ does not divide $(p_{0}-1)q_{0}$, $T_{1}(p_{0}-1,n(p_{0}-1)q_{0}/p_{0})$ is not the $n$-sheeted cover of $T_{1}(p_{1},q_{1})$. Hence, in any case, we have $l_{0}=l_{1}$, $p_{0}=p_{1}$, and $q_{0}=q_{1}$ by Lemma 3.1. Theorefore $L_{0}$ and $L_{1}$ are isotopic. ∎
@@
[…]
Suppose that $N>1$ and the assertion holds for any degree $n$ of the covering map if the number of JSJ pieces is less than $N$. Let $M_{0}$ and $M_{1}$ respectively denote the outermost JSJ pieces of $X\setminus\!\!\setminus L_{0}$ and $X\setminus\!\!\setminus L_{1}$. Their preimages $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ in $\widetilde{X}$ are isotopic in $\widetilde{X}$. The 3-manifolds $M_{0}$ and $M_{1}$ can be re-embedded as the complements of links in $X$ by Lemma 3.6. Here a knotted hole ball for $M_{i}$ is lifted to its disjoint $n$ copies in $\widetilde{X}$. The 3-manifolds $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ are equivariantly re-embedded. Since $M_{0}$ and $M_{1}$ are either Seifert fibered or hyperbolic, $M_{0}$ and $M_{1}$ are isotopic in $X$ by Theorem 4.4 and Lemma 4.6. Note that if $M_{i}$ is Seifert fibered and has a boundary component bounding a knotted hole ball, then $M_{i}$ is re-embedded as $X\setminus\!\!\setminus T_{1}(0,q)$.
@@
[…]
@@
Let $M_{0}$ and $M_{1}$ respectively denote the outermost JSJ pieces of $X\setminus\!\!\setminus L_{0}$ and $X\setminus\!\!\setminus L_{1}$ as in Lemma 3.12. Their preimages $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ in $\widetilde{X}$ are isotopic in $\widetilde{X}$. In the same manner as Theorem 4.5, $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ are equivariantly re-embedded in the thickened torus. Since $M_{0}$ and $M_{1}$ are either Seifert fibered or hyperbolic, $M_{0}$ and $M_{1}$ are isotopic in $X$ by Theorem 4.4 and Lemma 4.8.
@@
[…]